OpenStudy (anonymous):

Determine the slope between the points (-1, -5) and (5, 4) A. 2/3 B. -1/4 C. 1/4 D. 3/2

hero (hero):

y = mx + b -5 = -1m + b 4 = 5m + b 1m - 5 = b 4 - 5m = b b = b m - 5 = 4 - 5m m + 5m = 5 + 4 6m = 9 m = 9/6 m = 3/2

OpenStudy (anonymous):

\[slope = \frac{\Delta y}{\Delta x}\] \[\Delta y=4-(-5)\] \[\Delta x=5-(-1)\] slope=9/6=3/2